OpenStudy (anonymous):

Let f(x) = x^2-5x+2 (a) Find and simplify the difference quotient f(x+h)-f(x)/h Find f'(x) by evaluating lim h-> 0 f(x+h)-f(x)/h using the results from part (a).

OpenStudy (jamesj):

Ok, so write out \[ \frac{f(x+h) - f(x)}{h} \] and simplify. what do you get?

OpenStudy (anonymous):

No idea how to do step 1

OpenStudy (dumbcow):

\[f(x+h) = (x+h)^{2} -5(x+h)+2\] the variable changes from "x" to "x+h"

OpenStudy (jamesj):

First of all write out f(x+h) - f(x) and simplify f(x+h) - f(x) = (x+h)^2 - 5(x+h) + 2 - (x^2 - 5x +2) = ....